OpenStudy (vera_ewing):

A square has side lengths 12 in. The perimeter is 48 in cm and the area is 144 in2. A second square has side lengths that are one and a half times the first square. Find the area of the second square. A. 169 in2 B. 324 in2 C. 180 in2 D. 144 in2

OpenStudy (vera_ewing):

@TheRealMeeeee please help

OpenStudy (therealmeeeee):

@vera_ewing sorry I can't I forgot how to do this stuff,but I may know someone who can help you!! xD

OpenStudy (vera_ewing):

i've been waiting so long and i haven't gotten any help. @mathslover please help

OpenStudy (therealmeeeee):

@mathmale

OpenStudy (therealmeeeee):

@vera_ewing he is here the Best Math person in the world will help you

mathslover (mathslover):

@vera_ewing - Sorry I was away from keyboard. I will try to help you.

OpenStudy (vera_ewing):

@mathslover oh thank you!!! :))

OpenStudy (mathmale):

Hello, Vera! Looks as though you suddenly have gone from 0 helpers to 2 helpers! I'm going to let mathslover help with this problem, but will keep an eye on the progress you two make and step in if need be. @mathslover: thanks for offering help!

OpenStudy (vera_ewing):

@mathmale thank you!!! :)

mathslover (mathslover):

It is given to you that the sides of the second square are 1 and half times the sides of first square. Also, it is given that side of first square = 12 in. So, side of second square = \(1\cfrac{1}{2} \times \text{Side of First square}\)

OpenStudy (vera_ewing):

how would i do that?

mathslover (mathslover):

See, let us first convert \(1\cfrac{1}{2}\) (mixed fraction) to a simple fraction

OpenStudy (vera_ewing):

ok...

mathslover (mathslover):

It can be written as : \(\cfrac{2\times 1 + 1}{2}\) = \(\cfrac{3}{2}\)

mathslover (mathslover):

\(1\cfrac{1}{2}\) actually means : 1 (whole) + 1/2 So what I did was I just added these and got 3/2

OpenStudy (vera_ewing):

ok

OpenStudy (vera_ewing):

then what?

mathslover (mathslover):

Now , see, we have : side of second square = \(\cfrac{3}{2}\) * Side of First square it is given that side of first square is 12 in. So, side of second square = \(\cfrac{3}{2}\) * 12 in.

mathslover (mathslover):

Can you solve that and find the length of side of second square?

OpenStudy (vera_ewing):

no...i don't know how :(

OpenStudy (vera_ewing):

i can't multiply fractions

mathslover (mathslover):

Okay, no problem, I will teach you that. See, we have : \(\text{Side of Second Square} = \cfrac{3}{2} \times 12 \) Can you tell me what is \(\cfrac{12}{2}\) ?

OpenStudy (vera_ewing):

6

mathslover (mathslover):

Good! So, I can write that as : Side of Second Square = \(3 \times \cfrac{12}{2} = 3 \times 6\) Right?

OpenStudy (vera_ewing):

yes

mathslover (mathslover):

Okay, so now , what is 6 * 3?

OpenStudy (vera_ewing):

18

mathslover (mathslover):

Very Good! So, side of second square = 18 inches.

OpenStudy (vera_ewing):

ok...and then i multiply 18 by what?

mathslover (mathslover):

We have covered the first milestone now :) Now, it's the turn for the second and final milestone! Do you know what is the formula for Area of Square?

OpenStudy (vera_ewing):

LxW

mathslover (mathslover):

And since, L and W for a square are same, so it becomes (side)^2 ... okay ?

OpenStudy (vera_ewing):

OK, so the answer is 36?

OpenStudy (vera_ewing):

169 in2 324 in2 180 in2 144 in2

OpenStudy (vera_ewing):

^Those are my choices

mathslover (mathslover):

NO! \((side)^2 = side \times side\)

OpenStudy (vera_ewing):

oh 324^2

mathslover (mathslover):

It is 324 but now ^2 .. You just need to add unit to that. 324 square inches

OpenStudy (vera_ewing):

ok thanks!

OpenStudy (mathmale):

You two have worked together well, and I'm very happy about that.

OpenStudy (therealmeeeee):

Good Job @vera_ewing and @mathslover y'all make a wonderful team xD

mathslover (mathslover):

You're welcome. And thanks to @mathmale for giving me a chance to help you! Sorry for late help. And good luck!